3.
2(x + 3)2 = 50
If you were solving this equation by the square root method, what would you write for the next step of the problem?



Answer by nerdybill(7384)

2(x + 3)^2 = 50
divide both sides by 2:
(x + 3)^2 = 25
take the square root of both sides:
x+3 = +-5
x = +-5 - 3
.
x = {-8, 2}